A trapezoid has an area of 493 meters squared, a height of 17m, and a base of 22m. Find the length of the other base. Show all work to receive credit.
Answer by stanbon(75887)

A trapezoid has an area of 493 meters squared, a height of 17m, and a base of 22m. Find the length of the other base. Show all work to receive credit.
--------------------
Area = (1/2)h(b1+b2)
493 = (1/2)*17(22 + x)
---
984 = 17(22+x)
22+x = 57.88
x = 35.88
==================
Cheers,
Stan H.
==================